# Density-based clustering algorithm for galaxy group/cluster identification

### v1

@@ 3.1 Galaxy Sample
The Millennium Simulation tracks the evolution of $N=2160^{3}$ dark matter particles within a comoving volume of $500\,\mathrm{h}^{-1}\mathrm{Mpc}$ using the N-body code GADGET-2 (Springel, 2005). Sixty-four snapshots were periodically saved, along with group catalogs and their substructures identified through a two-step process. First, the FoF algorithm with a linking length of 0.2 in units of the mean particle separation identified potential halos. These candidates were then refined by the SUBFIND algorithm (Springel et al., 2001) through a gravitational unbinding procedure, ensuring only substructures with at least 20 particles were considered to be genuine halos and substructures. Subsequently, with halos detailed merger history trees were constructed for all gravitationally bound structures in each snapshot. The merger trees trace the evolution of these structures throughout cosmic time, providing the crucial temporal and structural framework upon which SAMs operate. Within this framework, SAMs simulate the formation and evolution of galaxies, ultimately populating the dark matter halos with galaxies. (De Lucia & Blaizot, 2007)
@@ 3.2 Comparing Clustering Algorithms
Considering the hierarchical structure of the Universe, with galaxy groups typically hosting 3 to 30 bright galaxies and clusters holding 30 to over 300, it is logical to focus our search for optimal clustering parameters within this range. The Local Group, for instance, hosts over 30 galaxies with a diameter of nearly 3 Mpc (McConnachie et al., 2005). Therefore, we set the searching space of distance thresholds for clustering algorithms between 0.1 Mpc and a few Mpc. Similarly, the minimum neighbor number $N_{\mathrm{min}}$ and minimum member number $M_{\mathrm{min}}$ are explored within the range of 2-20. We employ a broader range, 500 to 5000, for algorithms requiring a preselected number of clusters, to ensure exploring all possibilities. The trial hyperparameter values for all algorithms are listed in Table 1.
@@
Friend-of-friends linking_length $l$ [0.001, 3.0) 0.18 83.5% 1114
max_eps $\epsilon$ [0.05, 1.0) 0.2
OPTICS min_sample $N_{\mathrm{min}}$ [2, 20) 3 83.0% 1035
@@
min_member $M_{\mathrm{min}}$ [2, 20) 3
DBSCAN max_eps $\epsilon$ [0.05, 1.0) 0.2
min_sample $N_{\mathrm{min}}$ [2, 20) 2 84.0% 2778
min_member $M_{\mathrm{min}}$ [2, 20) 2
@@
HDBSCAN min_member $M_{\mathrm{min}}$ [2, 20) 2 54.5% 3433
@@
As we can see from the results, the traditional methods FoF, OPTICS, and DBSCAN can effectively recover the galaxy groups just based on the spatial distribution of galaxies with a recovery rate of over 70%. The others can not give a good prediction of the groups. It should also be emphasized that the ground was calculated based on the FoF algorithm.
@@ 3.3 Parameter Sensitivity
Even though FoF and OPTICS are comparable in predicting galaxy groups, they differ significantly in their hyperparameter complexities. OPTICS requires tuning four hyperparameters, providing more flexibility and necessitating more careful configuration. On the other hand, FoF has only one primary hyperparameter, simplifying its use but potentially limiting its adaptability. This contrast raises questions about the sensitivity of their respective hyperparameters. To investigate this, purity, completeness, and recovery rates were calculated for each algorithm under different values of a single hyperparameter while keeping the others constant. Figure 1 presents results based on soft criteria. These analyses were conducted on the same subsample described in Section 3.1.
@@ 4.2 Cure the Redshift-Space Distortion via sOPTICS with a LOS Scaling Factor
To address redshift-space distortion, we introduce a LOS scaling factor denoted as $s_{\mathrm{LOS}}$ for the line-of-sight component. Consequently, the modified distance metric, referred to as the Elongated Euclidean Distance, is computed as follows:
$D_{\mathrm{Elongated}}^{2}(u,v,s_{\mathrm{LOS}})=d_{\mathrm{Transverse}}^{2}(u,v)+d_{\mathrm{LOS,\,scaled}}^{2}(u,v,s_{\mathrm{LOS}})\;.$ (5)
The LOS component is calculated by projecting the difference vector:
$d_{\mathrm{LOS}}(u,v)=\frac{\sum_{i=1}^{3}(u_{i}-v_{i})u_{i}}{\sqrt{\sum_{i=1}^{n}u_{i}^{2}}}\;.$ (6)
Subsequently, this component is scaled by $s_{\mathrm{LOS}}$, representing the LOS scaling factor, to obtain the scaled LOS component:
$d_{\mathrm{LOS,scaled}}(u,v,s_{\mathrm{LOS}})=s_{\mathrm{LOS}}\,d_{\mathrm{LOS}}(u,v)\;.$ (7)
The final distance metric, the elongated Euclidean distance, incorporates this scaled LOS component, effectively compensating for the redshift-space distortion by "shortening" the distance along the line of sight.
@@
The approximation $z=z_{\rm c}+z_{\mathrm{p}}$ is only valid for small redshifts. Consequently, redshift-space distortion increasingly affects galaxy groups at higher redshifts. This stronger distortion necessitates more robust adjustments, specifically, smaller $s_{\mathrm{LOS}}$. Therefore, it is necessary to adjust the value of the LOS scaling factor $s_{\mathrm{LOS}}$ with redshift. To ascertain the proper values of $s_{\mathrm{LOS}}$ correcting for redshift-space distortions across varying redshift bins, in Section 4.3 we iteratively adjusted $z=z_{\rm c}+z_{\mathrm{p}}$ to optimize the concordance between our observations and the Y07 group catalog. Figure 5 illustrates how the optimal LOS scaling factor changes with redshift. It is clearly shown that the optimal LOS scaling factor decreases with higher redshifts, indicating that the Euclidean distance along the LOS is elongated more significantly. This trend is consistent with theoretical predictions.
@@ 4.3 Choices of Hyperparameter Values
To fine-tune the hyperparameters, we select ten subsamples from low redshift galaxies ($z<0.05$), each with a cubic side length of 100 Mpc. Then, we first identify the optimal values of hyperparameters of FoF, as well as sOPTICS with a constant $s_{\mathrm{LOS}}$, which maximizes the abundance-weighted recovery rate, as detailed in Table 2. Using these hyperparameters, we achieved a maximum recovery rate of $F_{\mathrm{R}}=0.8$ for FoF and 0.76 for sOPTICS. It is important to note that for sOPTICS, although the recovery rate of 0.76 wasn’t the peak for every individual test subsample—with the highest rate reaching 0.89 in certain scenarios—these hyperparameters yield the most consistent and accurate predictions of BCGs across the board, as elaborated in Section 5.1. Thus, we adopted this set of hyperparameters for sOPTICS as the most suitable choice, balancing overall performance across various testing conditions.
@@ 4.4 Basic Results
With the optimal values for hyperparameters and the LOS scaling factor listed in Table 2 and 3, the overall abundance-weighted recovery rate for galaxy groups using the sOPTICS algorithm is 74.9%, with the abundance-weighted purity of 86.6% and completeness of 96.9%. The total number of identified groups is 12,196. Meanwhile, the soft recovery rate stands at 69.0%, indicating that we can precisely identify 5,811 galaxy groups, with two-thirds of their member galaxies matching those of the true groups identified in Y07 and covering more than half of the actual members, out of a total of 8,427 true groups. In contrast, the FoF algorithm achieved a soft recovery rate of 64.7%, while with a purity of 46.3% and completeness of 78.9%. The total number of identified groups using FoF is 95,732, which is much larger than the number identified by sOPTICS, indicating a very low efficiency identification and non-negligible contamination in the detected group. Therefore, by incorporating the LOS scaling factor, we significantly enhanced the precision of the sOPTICS algorithm’s predictions. Consequently, in identifying galaxy groups and clusters, sOPTICS performs comparably to, and in some cases slightly better than, FoF when parameters are tuned based on sub-samples.
@@ 5.1 Performance of finding BCGs
As demonstrated, sOPTICS can effectively detect large clusters with precise member identification, including the BCGs. To evaluate the performance of our sOPTICS method in identifying BCGs from a galaxy survey, we conducted a comparative test against a recent BCGs sample from Hsu et al. (2022, hereafter Hsu22). Their parent BCG sample of 4,033 galaxy clusters is also extracted from the group catalog of Y07 with applying a cut in the cluster mass $M_{180m}\geqslant 10^{14}h^{-1}M_{\odot}$. By cross-matching BCG candidates with the 8,113 galaxies released in the ninth Product Launch (MPL-9) of the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA Bundy et al., 2014), they identified 128 BCGs situated within a redshift range of $z=0.02-0.15$. These clusters are all detected in the X-rays by Wang et al. (2014), which provides a cluster catalog with X-ray luminosity from ROSAT All Sky Survey. However, Y07 primarily select BCGs based on luminosity, occasionally resulting in the selection of spiral galaxies as BCG candidates. Therefore, Hsu22 implemented an additional visual selection process: if the BCG candidates show a spiral morphology or do not represent the most luminous galaxy on the red sequence, alternative candidates would considered. The cluster would be excluded if no superior candidate exists or MaNGA has not observed the more suitable candidate. As a result, 121 BCGs have been visually confirmed, of which 118 were originally part of the Y07 catalog.
@@
Using the best-fitted parameters and the LOS scaling factor listed in Table 2 and 3, we successfully identified 115 BCGs consistent with the 118 BCGs identified in Hsu22. The spatial distribution of the galaxy clusters corresponding to these BCGs in redshift and right ascension (RA) space is illustrated in Figure 7. Only three relatively small clusters failed to be predicted by sOPTICS. Figure 8 shows a segment of the reachability plot for scaled OPTICS, where the galaxy clusters appear as distinct, deep valleys. The gray areas represent isolated field galaxies that are significantly distanced from others. The bottom panel shows a specific example cluster’s reachability distances and neighbors, including a BCG recovered from the Hsu22 sample. It is shown that these density-based clustering methods, such as OPTICS and sOPTICS, have given us a clear and straightforward picture of the position of BCGs in clusters. In this particular case, the BCG is located precisely at the densest part of the region, indicating a perfect alignment with the cluster’s center of gravity. However, it is also evident from other commonly detected cases (highlighted as yellow spots in the plot) that the BCGs are not always situated at the densest part of the cluster. This variation highlights the diversity in the spatial distributions of BCGs and explains why our sOPTICS method did not successfully predict three BCGs.

### v2

@@ 3.1 Galaxy Sample
The Millennium Simulation tracks the evolution of $N=2160^{3}$ dark matter particles within a comoving volume of $500\,\mathrm{h}^{-1}\mathrm{Mpc}$ using the N-body code GADGET-2 (Springel, 2005). Sixty-four snapshots were periodically saved, along with group catalogs and their substructures identified through a two-step process. First, the FoF algorithm with a linking length of 0.2 in units of the mean particle separation identified potential halos. These candidates were then refined by the SUBFIND algorithm (Springel et al., 2001) through a gravitational unbinding procedure, ensuring only substructures with at least 20 particles were considered to be genuine halos and substructures. Subsequently, with halos detailed merger history trees were constructed for all gravitationally bound structures in each snapshot. The merger trees trace the evolution of these structures throughout cosmic time, providing the crucial temporal and structural framework upon which SAMs operate. Within this framework, SAMs simulate the formation and evolution of galaxies, ultimately populating the dark matter halos with galaxies (De Lucia & Blaizot, 2007).
@@ 3.2 Comparing Clustering Algorithms
Considering the hierarchical structure of the Universe, with galaxy groups typically hosting 3 to 30 bright galaxies and clusters holding 30 to over 300, it is logical to focus our search for optimal clustering parameters within this range. The Local Group, for instance, hosts over 30 galaxies with a diameter of nearly 3 Mpc (McConnachie et al., 2005). Therefore, we set the upper limit for the linking length $l$ in FoF, $\epsilon$ in OPTICS and DBSCAN to 3.0 Mpc, corresponding to a reasonable maximum for the spatial sparsity of galaxies within a group. Similarly, the minimum neighbor number $N_{\mathrm{min}}$ and minimum member number $M_{\mathrm{min}}$ are explored within the range of 2-20. We employ a broader range, 500 to 5000, for algorithms requiring a preselected number of clusters, to ensure exploring all possibilities. The trial hyperparameter values for all algorithms are listed in Table 1.
@@
Friend-of-friends linking_length $l$ [0.001, 3.0) 0.24 79.0% 471
max_eps $\epsilon$ [0.05, 3.0) 0.25
OPTICS min_sample $N_{\mathrm{min}}$ [2, 20) 3 79.2% 472
@@
min_member $M_{\mathrm{min}}$ [5, 20) 5
DBSCAN max_eps $\epsilon$ [0.05, 3.0) 0.2
min_sample $N_{\mathrm{min}}$ [2, 20) 2 75.2% 406
min_member $M_{\mathrm{min}}$ [5, 20) 5
@@
HDBSCAN min_member $M_{\mathrm{min}}$ [5, 20) 5 49.5% 1576
@@
As we can see from the results, the traditional methods FoF, OPTICS, and DBSCAN can effectively recover the galaxy groups just based on the spatial distribution of galaxies with a recovery rate of over 70%. The others, including HDBSCAN, can not give a good prediction of the groups. It should also be emphasized that the ground truth was calculated based on the FoF algorithm.
In particular, $k$-means, GMMs, Agglomerative Clustering, and Spectral Clustering all demonstrate significant shortcomings in this context. Their results clearly show that they are not suitable for identifying galaxy clusters. This is due to the nature of the algorithms themselves: they assume fixed or simplistic cluster shapes (e.g., spherical or Gaussian) and rely on predefined scales, which do not align with the complex, hierarchical, and irregular distribution of galaxies in cosmic structures. What is worse is that Spectral Clustering not only struggles with the complexity of the task but is also computationally expensive. Even for a moderately sized sample much smaller than real observational datasets, the algorithm required excessive runtime to complete, rendering it impractical for larger-scale applications. Consequently, we conclude that these four algorithms are unsuitable for finding galaxy clusters and will not be included in the following tests or discussions.
@@ 3.3 Parameter Sensitivity
Even though FoF and OPTICS are comparable in predicting galaxy groups, they differ significantly in their hyperparameter complexities. OPTICS requires tuning four hyperparameters, providing more flexibility and necessitating more careful configuration. On the other hand, FoF has only one primary hyperparameter, simplifying its use but potentially limiting its adaptability. This contrast raises questions about the sensitivity of their respective hyperparameters. To investigate this, purity, completeness, and recovery rates were calculated for each algorithm under different values of a single hyperparameter while keeping the others unchanged at their respective optimal values. Figure 1 presents results based on soft criteria. These analyses were conducted on the same subsample described in Section 3.1.
@@ 4.2 Cure the Redshift-Space Distortion via sOPTICS with a LOS Scaling Factor
To address redshift-space distortion, we introduce a LOS scaling factor denoted as $s_{\mathrm{LOS}}$ for the line-of-sight component. This factor is applied specifically to the LOS component, thereby “shortening” distances in that direction and mitigating the effects of redshift-space distortion. We define the Elongated Euclidean Distance as:
$D_{\mathrm{Elongated}}^{2}\bigl{(}u,v,s_{\mathrm{LOS}}\bigr{)}=d_{\mathrm{Transverse}}^{2}(u,v)+d_{\mathrm{LOS,,scaled}}^{2}\bigl{(}u,v,s_{\mathrm{LOS}}\bigr{)}.$ (6)
Here, $d_{\mathrm{Transverse}}(u,v)$ corresponds to the component of the Euclidean distance perpendicular to the LOS. And $d_{\mathrm{LOS,\,scaled}}$ is the scaled LOS component. In principle, one could define:
$d_{\mathrm{LOS}}(u,v)=\frac{\sum_{i=1}^{3}\bigl{(}u_{i}-v_{i}\bigr{)},u_{i}}{\sqrt{\sum_{i=1}^{3}\bigl{(}u_{i}\bigr{)}^{2}}},,$ (7)
and then multiply by $s_{\mathrm{LOS}}$ to obtain:
$d_{\mathrm{LOS,,scaled}}\bigl{(}u,v,s_{\mathrm{LOS}}\bigr{)}=s_{\mathrm{LOS}}d_{\mathrm{LOS}}(u,v).$ (8)
However, since $d_{\mathrm{LOS}}(u,v)$ does not equal $d_{\mathrm{LOS}}(v,u)$, the resulting distance may lose symmetry. In other words, $D_{\mathrm{Elongated}}(u,v)\neq D_{\mathrm{Elongated}}(v,u)$, which can undermine the metric properties typically assumed by OPTICS and potentially increase the algorithm’s sensitivity to data ordering. To ensure a symmetric distance measure and preserve the stability of core-point definitions, we adopt a symmetrized version of the LOS component:
$d_{\mathrm{LOS,,scaled}}^{\mathrm{sym}}\bigl{(}u,v,s_{\mathrm{LOS}}\bigr{)}=s_{\mathrm{LOS}}\frac{d_{\mathrm{LOS}}(u,v)+d_{\mathrm{LOS}}(v,u)}{2}.$ (9)
This modification guarantees that $d_{\mathrm{LOS,\,scaled}}^{\mathrm{sym}}(u,v,s_{\mathrm{LOS}})=d_{\mathrm{LOS,\,scaled}}^{\mathrm{sym}}(v,u,s_{\mathrm{LOS}})$, thereby restoring symmetry to the overall distance function. By enforcing such symmetry, one retains the theoretical benefits of a core-distance-based reachability measure in OPTICS—namely, stable cluster structures that are less dependent on the processing order of data points—and ensures that the clustering results remain interpretable as a function of an actual distance metric.
@@
The approximation $z=z_{\rm c}+z_{\mathrm{p}}$ is only valid for small redshifts. Consequently, redshift-space distortion increasingly affects galaxy groups at higher redshifts. This stronger distortion necessitates more robust adjustments, specifically, smaller $s_{\mathrm{LOS}}$. Therefore, it is necessary to adjust the value of the LOS scaling factor $s_{\mathrm{LOS}}$ with redshift. To ascertain the proper values of $s_{\mathrm{LOS}}$ correcting for redshift-space distortions across varying redshift bins, in Section 4.3 we iteratively optimized $s_{\mathrm{LOS}}$ by maximizing the concordance between our results and the Y07 group catalog. Figure 5 illustrates how the optimal LOS scaling factor changes with redshift. It is clearly shown that the optimal LOS scaling factor decreases with higher redshifts, indicating that the Euclidean distance along the LOS is elongated more significantly. This trend is consistent with theoretical predictions.
@@ 4.3 Choices of Hyperparameter Values
To fine-tune the hyperparameters, we select ten subsamples from low redshift galaxies ($z<0.05$), each with a cubic side length of 100 Mpc. Then, we first identify the optimal values of hyperparameters of FoF, as well as sOPTICS with a constant $s_{\mathrm{LOS}}$. The search spaces for these hyperparameters are identical to that listed in Table 1, and the optimal values are detailed in Table 2. Using these hyperparameters, we achieved a maximum recovery rate of $F_{\mathrm{R}}=0.8$ for FoF and 0.76 for sOPTICS. It is important to note that for sOPTICS, although the recovery rate of 0.76 wasn’t the peak for every individual test subsample—with the highest rate reaching 0.89 in certain scenarios—these hyperparameters yield the most consistent and accurate predictions of BCGs across the board, as elaborated in Section 5.3. Thus, we adopted this set of hyperparameters for sOPTICS as the most suitable choice, balancing overall performance across various testing conditions.
@@ 4.4 Basic Results
With the optimal values for hyperparameters and the LOS scaling factor listed in Table 2 and 3, the overall abundance-weighted recovery rate for galaxy groups using the sOPTICS algorithm is 75.0%, with the abundance-weighted purity of 86.6% and completeness of 97.1%. The total number of identified groups is 12,242, while there are 8,427 groups with at least 5 member galaxies in Y07 catalog. Meanwhile, the soft recovery rate stands at 69.2%, indicating that we can precisely identify 5,831 galaxy groups, with two-thirds of their member galaxies matching those of the true groups identified in Y07 and covering more than half of the actual members, out of a total of 8,427 true groups. In contrast, the FoF algorithm achieved a soft recovery rate of 42.6%, while with a purity of 44.2% and completeness of 54.7% The total number of identified groups using FoF is 12,396. Therefore, by incorporating the LOS scaling factor, we significantly enhanced the precision of the sOPTICS algorithm’s predictions. Consequently, in identifying galaxy groups and clusters, sOPTICS performs comparably to, and in some cases even better than, FoF when parameters are tuned based on sub-samples.
@@ 5 Capability of Clustering Algorithms
5.1 Interdependence of Hyperparameters in sOPTICS
In employing the scaled OPTICS clustering algorithm to identify galaxy clusters, the hyperparameters $\epsilon$ (the maximum radius for neighborhood density estimation), $N_{\mathrm{min}}$ (the minimum number of points required to form a cluster), and the LOS scaling factor crucially influence the results, as detailed in Section 3. These parameters are pivotal in determining reachability distances and adjusting the algorithm to mitigate the effects of redshift-space distortion. Although one might anticipate a high sensitivity to parameter variations, sOPTICS exhibits resilience by maintaining an optimal range for these values. This finding is illustrated in Figures 7 and 8, where the interdependence of $\epsilon$ and $N_{\mathrm{min}}$, as well as $\epsilon$ and $s_{\mathrm{LOS}}$, is presented. Notably, a clear correlation emerges between $\epsilon$ and $N_{\mathrm{min}}$; as $\epsilon$ increases, $N_{\mathrm{min}}$ must also to be adjusted upward to maintain effective clustering. Essentially, to preserve the purity of the clusters identified by the sOPTICS algorithm, the criteria must shift toward identifying denser and larger clusters as the $\epsilon$ threshold is raised. The positive correlation suggests that $\epsilon^{3}\propto N_{\mathrm{min}}$ aligns with theoretical expectations. Theoretically, this adjustment ensures that the increase in the neighborhood radius does not lead to the inclusion of outlier points or less dense areas. Thus $N_{\mathrm{min}}$ should correlate with the volume of space encompassed within $\epsilon$, which implies a cubic relationship ($\epsilon^{3}$).
Figure 7: The results of galaxy group finding for various pairs of values of $\epsilon$ and $N_{\mathrm{min}}$. The color bar represents the weighted recovery rate of the predicted groups relative to a subsample from the Y07 groups catalog, covering a $30^{2}\,\deg^{2}$ sky area within the redshift range $0.80.7<z<$ 0.8.
Figure 8: As Fig. 7, but showing the results of galaxy group finding for various pairs of values of $s_{\mathrm{LOS}}$ and $\epsilon$.
These relationships underscore the interconnected nature of $\epsilon$, $N_{\mathrm{min}}$, and $s_{\mathrm{LOS}}$, as well as their collective impact on optimizing cluster detection and recovery rates. Specifically, selecting parameter values within this optimal range can yield galaxy groups with a high recovery rate comparable to those obtained from reliable group catalogs that require complex and computationally intensive processes. The optimal ranges depicted in Figures 7 and 8 also identify a potential characteristic number density for categorizing galaxies in a survey as a group. Given the similar local completeness of the survey, this characteristic number density of galaxy groups can be applied to other observations.
However, given the current redshift range of the observed data, the relationships observed between $\epsilon$, $N_{\mathrm{min}}$, and $s_{\mathrm{LOS}}$ are preliminary and roughly empirical. To more precisely define these relationships and understand the characteristic number density, a comprehensive analysis using both real-world data and mock catalogs is crucial. This approach would help determine whether the observed linear trend between $\epsilon$ and $s_{\mathrm{LOS}}$ is an artifact of the specific dataset used in this study or if it reflects a more general characteristic applicable across different galaxy cluster distributions.
5.2 Application of sOPTICS to an Independent Catalog
Beyond examining empirical parameter relationships in a controlled setting, a core advantage of sOPTICS is its adaptability to other observational datasets with minimal re-tuning of parameters. To explore this aspect, we applied sOPTICS to an independent catalog from Shi et al. (2016), which contains 586,025 galaxies from SDSS DR13 (Albareti et al., 2017) spanning a redshift range of $0\leq z\leq 0.20$. The group catalog was constructed using the adaptive halo-based group finder developed by Yang et al. (2005); Yang et al. (2007). Specifically, they corrected redshift-space distortions by reconstructing the large-scale velocity field to account for the Kaiser effect and statistically redistributing galaxies within halos based on an NFW profile to address the Finger of God effect, yielding a pseudo-real-space galaxy group catalog of 8,640 systems (each containing at least five member galaxies).
We adopt the same sOPTICS hyperparameters determined in Tables 2 and 3, without additional fine-tuning. Under these settings, sOPTICS identified 10,057 predicted groups, achieving a recovery rate of 55.2%, a purity of 63.7%, and a completeness of 75.8%. These results are notable because they demonstrate that sOPTICS’s parameter choices, optimized on a different dataset, remain effective when transferred to new observations.
In contrast, we apply the FoF linking length obtained from Tables 2 to the same catalog. This yield significantly lower performance: 9,421 predicted groups with a recovery rate of 30.8%, a purity of 46.7%, and a completeness of 50.2%. The substantial disparity suggests that FoF is more sensitive to its linking length parameter, whereas sOPTICS is comparatively robust to modest changes in data characteristics. By adapting cluster boundaries based on local density structures, sOPTICS retains its effectiveness across multiple samples, reducing the need for exhaustive parameter searches.
To investigate whether performance could be enhanced further, we examined the interdependence of $\epsilon$ and $N_{\mathrm{min}}$ on this new dataset while keeping other sOPTICS parameters unchanged (as in Figure 7). We find the same correlation pattern between these two parameters that was observed in our earlier tests, indicating that increasing $\epsilon$ necessitates higher values of $N_{\mathrm{min}}$ to preserve cluster purity. By simply optimizing $\epsilon$ and $N_{\mathrm{min}}$ alone, the recovery rate improved to 67%, as presented in Figure 9. This outcome underscores both the flexibility of sOPTICS and its potential for achieving higher cluster identification accuracy through modest parameter adjustments.
Figure 9: As Fig. 7, but showing the validation results of the predicted groups relative to a subsample from the groups catalog of Shi et al. (2016).
@@ 5.3 Performance of finding BCGs
As demonstrated, sOPTICS can effectively detect large clusters with precise member identification, including the BCGs. To evaluate the performance of our sOPTICS method in identifying BCGs from a galaxy survey, we conducted a comparative test against a recent BCGs sample from Hsu et al. (2022, hereafter Hsu22). Their parent BCG sample of 4,033 galaxy clusters is also extracted from the group catalog of Y07 with applying a cut in the cluster mass $M_{180m}\geqslant 10^{14}h^{-1}M_{\odot}$. By cross-matching BCG candidates with the 8,113 galaxies released in the ninth Product Launch (MPL-9) of the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA Bundy et al., 2014), they identified 128 BCGs situated within a redshift range of $z=0.02-0.15$. These clusters are all detected in the X-rays by Wang et al. (2014), which provides a cluster catalog with X-ray luminosity from the ROSAT All Sky Survey. However, Y07 primarily select BCGs based on luminosity, occasionally resulting in the selection of spiral galaxies as BCG candidates. Therefore, Hsu22 implemented an additional visual selection process: if the BCG candidates show a spiral morphology or do not represent the most luminous galaxy on the red sequence, alternative candidates would considered. The cluster would be excluded if no superior candidate exists or MaNGA has not observed the more suitable candidate. As a result, 121 BCGs have been visually confirmed, of which 118 were originally part of the Y07 catalog.
@@
Using the best-fitted parameters and the LOS scaling factor listed in Table 2 and 3, we successfully identified 115 BCGs consistent with the 118 BCGs identified in Hsu22. The spatial distribution of the galaxy clusters corresponding to these BCGs in redshift and right ascension (RA) space is illustrated in Figure 10. Only three relatively small clusters failed to be predicted by sOPTICS. Figure 11 shows a segment of the reachability plot for sOPTICS, where the galaxy clusters appear as distinct, deep valleys. The gray areas represent isolated field galaxies that are significantly distanced from others. The bottom panel shows a specific example cluster’s reachability distances and neighbors, including a BCG recovered from the Hsu22 sample. It is shown that these density-based clustering methods, such as OPTICS and sOPTICS, have given us a clear and straightforward picture of the position of BCGs in clusters. In this particular case, the BCG is located precisely at the densest part of the region, indicating a perfect alignment with the cluster’s center of gravity. However, it is also evident from other commonly detected cases (highlighted as yellow spots in the plot) that the BCGs are not always situated at the densest part of the cluster. This variation highlights the diversity in the spatial distributions of BCGs and explains why our sOPTICS method did not successfully predict three BCGs.